40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 16 23:30:34 2021, Paco, Update, CDS, AS WFS commissioning; restarting models 
    Reply  Thu Aug 19 03:23:00 2021, Anchal, Update, CDS, Time synchornization not running 
       Reply  Thu Aug 19 14:14:49 2021, Koji, Update, CDS, Time synchornization not running 
          Reply  Fri Aug 20 00:28:55 2021, Anchal, Update, CDS, Time synchornization not running 
             Reply  Fri Aug 20 06:24:18 2021, Anchal, Update, CDS, Time synchornization not running 
Message ID: 16280     Entry time: Mon Aug 16 23:30:34 2021     Reply to this: 16283
Author: Paco 
Type: Update 
Category: CDS 
Subject: AS WFS commissioning; restarting models 

[koji, ian, tega, paco]

With the remote/local assistance of Tega/Ian last friday I made changes on the c1sus model by connecting the C1:ASC model outputs (found within a block in c1ioo) to the BS and PRM suspension inputs (pitch and yaw). Then, Koji reviewed these changes today and made me notice that no changes are actually needed since the blocks were already in place, connected in the right ports, but the model probably just wasn't rebuilt...

So, today we ran "rtcds make", "rtcds install" on the c1ioo and c1sus models (in that order) but the whole system crashed. We spent a great deal of time restarting the machines and their processes but we struggled quite a lot with setting up the right dates to match the GPS times. What seemed to work in the end was to follow the format of the date in the fb1 machine and try to match the timing to the sub-second level. This is especially tricky when performed by a human action so the whole task is tedious. We anyways completed the reboot for almost all the models except the c1oaf (which tends to make things crashy) since we won't need it right away for the tasks ahead. One potential annoying issue we found was in manually rebooting c1iscey because one of its network ports is loose (the ethernet cable won't click in place) and it appears to use this link to boot (!!) so for a while this machine just wasn't coming back up.

Finally, as we restored the suspension controls and reopened the shutters, we noticed a great deal of misalignment to the point no reflected beam was coming back to the RFPD table. So we spent some time verifying the PRM alignment and TT1 and TT2 (tip tilts) and it turned out to be mostly the latter pair that were responsible for it. We used the green beams to help optimize the XARM and YARM transmissions and were able to relock the arms. We ran ASS on them, and then aligned the PRM OpLevs which also seemed off. This was done by giving a pitch offset to the input PRM oplev beam path and then correcting for it downstream (before the qpd). We also adjusted the BS OpLev in the end.


Summary; the ASC BS and PRM outputs are now built into the SUS models. Let the AS WFS loops be closed soon!


Addenda by KA
- Upon the RTS restarting,

  • Date/Time adjustment
    sudo date --set='xxxxxx'
  • If the time on the CDS status medm screen for each IOP match with the FB local time, we ran
    rtcds start c1x01
    (or c1x02, etc)
  • Every time we restart the IOPs, fb was restarted by
    telnet fb1 8083
    > shutdown

    and restarted mx_stream from the CDS screen because these actions change the "DC" status.

- Today we once succeeded to restart the vertex machines. However, the RFM signal transmission did fail. So the end two machines were power cycled as well as c1rfm, but this made all the machines in RED again. Hell...

- We checked the PRM oplev. The spot was around the center but was clipped. This made us so confused. Our conclusion was that the oplev was like that before the RTS reboot.

ELOG V3.1.3-